# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ 3.6.3 Other Applications
Once the method for quantifying causal emergence in complex systems is developed, it can be applied across various fields that possess abundant network data. The first type of networks which was studied is biological networks.
@@
As previously discussed, biological networks are full of noise, which poses challenges in comprehending their internal operating principles. On the one hand, such noise arises from the inherent fluctuations within the system itself, while on the other hand, it can be introduced through measurement or observation processes. Consequently, Klein et al. Klein et al. (2022) further explored the relationship among noise, degeneracy and certainty in biological networks and their specific meanings. For instance, in gene expression networks, highly deterministic relationships indicate that the expression of one gene almost invariably leads to the expression of another gene. Simultaneously, degeneracy is a prevalent phenomenon in the evolutionary processes of biological systems. Due to these two factors, it remains unclear at which scale biological systems should be analyzed to gain a deeper understanding of their functions.
@@
To address this, Klein et al. Klein et al. (2021) conducted an analysis of protein interaction networks across more than 1,800 species. They employed EI as a measure to assess the levels of noise and uncertainty in protein interactions. The findings revealed that the macro-scale network exhibited lower levels of noise and degeneracy. Additionally, the nodes within the macro-scale interaction group demonstrated greater resilience compared to nodes that were not part of the macro-scale interaction group. Through robust analysis, the authors demonstrate that eukaryotes exhibit a stronger degree of causal emergence compared to archaea. Additionally, to address the ’deterministic paradox,’ the authors introduce the concepts of the neutral process and the selection process in biological evolution. The neutral process operates at the micro scale, leveraging mutations to promote interactions and enhance species diversity. On the other hand, the selection process operates at the macro scale, effectively eliminating noise that hampers system operation and efficiency. Therefore, in order to adapt to the demands of evolution, it becomes essential for evolved biological systems to function across multiple scales.
@@
Hoel et al. Hoel and Levin (2020) conducted further research on the causal emergence within biological systems. The authors elaborated that macro and micro systems exist widely in biological systems, for example, the micro scale of a group of cells can be the potential ion channel changes, and the macro scale corresponds to the membrane potential changes of cells. Furthermore, the authors utilized EI to analyze gene regulatory networks, aiming to identify the most informative models for controlling mammalian heart development. By quantifying the causal emergence within the largest component of the gene network of Saccharomyces cerevisia, it is revealed that informative macro-scale structures are prevalent across biological systems. Additionally, the authors emphasize the importance of evolutionary systems operating at multiple scales, as it offers significant advantages. Natural selection requires variation between populations, and degradation observed in biological systems serves as a crucial factor for evolution. Degradation provides the conditions for the evolutionary process. However, organisms also need to maintain predicted consistency in phenotype, behavior, and structure to ensure survival and reproduction. Consequently, evolved systems need to operate at multiple scales, and the function of multi-scale systems is malleable in changing environments.
@@
Swain et al. Swain et al. (2022) conducted an investigation into the impact of ant colony interaction history on task assignment and task switching. They employed effective information to examine how noise information propagates among ants, while also exploring the relationship between EI and the proportion of ants assigned to different tasks within ant colonies. The study revealed that the extent of historical interaction between ant colonies influenced task assignment. Additionally, the specific type of ant colony involved in an interaction determined the level of noise present within that interaction. For example, the interactions between foragers displayed significantly higher levels of noise when contrasted with interactions between nurses or cleaners. Furthermore, even when ants switched functional groups, the cohesion within ant colonies ensured the stability of the overall colony, and different functional ant colonies also play different roles in maintaining group cohesion.
@@
EI indicator and causal emergence theoretic framework can also be applied to artificial systems. For example, Marrow et al. Marrow et al. (2020) quantify and monitor the changes in the causal structure of neural networks during training, in which EI was employed to evaluate the degree of causal influence of nodes and edges on downstream tasks at each layer. By observing the changes of EI, including determinism(sensitivity) and degeneracy, throughout model training, the generalization ability of the model can be determined, thus helping us to better understand and explain the working principle of neural networks.
@@
Varley et al. Varley (2020) attempted to apply the causal emergence framework to both discrete cellular automata and continuous Rossler systems. In the case of cellular automata systems, the authors select 88 unique rules corresponding to four types: static, periodic, chaotic and complex. By considering each state as a node, where each state determines the subsequent state, a directed state transition graph is constructed. The analysis revealed that rules 1, 2, and 4 correspond to the strongest causal emergence of dynamics. Notably, the network constructed by these three rules exhibited a significant presence of star and spoke motifs. In addition, the authors also draw some quantitative conclusions, for example, among the 17 rules belonging to the third and fourth categories, 30% exist causal emergence, 70% show causal degradation, and the CE of cellular automata with the same rules remained relatively consistent across different sizes.
@@
Furthermore, the author employed the OPN algorithm to transform the continuous system into a discrete state transition graph for comparative analysis (see section 3.6.1). They found that chaos dynamics demonstrated a correlation with low determinism, and the variations in degeneracy and efficiency coefficients aligned with the changes observed in the determinism curve.
3.7 Critiques
@@
In the extensive literature on causality and emergence, Hoel’s theory has gained attention for its association of emergence and causality through interventionism, introducing the concept of causal emergence in a quantitative manner. However, Dewhurst  Dewhurst (2021) provides some philosophical clarity of Hoel’s theory, arguing that it is epistemological rather than metaphysical. It suggests that Hoel’s macroscopic causality is merely a causal explanation based on information theory, rather than involving "genuinely novel causal powers." This also raises concerns about the assumption of a uniform distribution, as there is no empirical basis to favor it over other distributions.
@@
The computation of Hoel’s effective information relies on two premises: 1) knowledge of the system’s microscopic dynamics, and 2) knowledge of the coarse-graining scheme. However, in practice, it is rare for both conditions to be simultaneously satisfied, especially in observational studies where both are unknown. This limitation hinders the practical applicability of Hoel’s theory.
@@
It is pointed out that Hoel’s theory negleted the constraints on coarse-graining methods, and certain coarse-graining methods can lead to ambiguity Eberhardt and Lee (2022). Additionally, the combination of some coarse-graining operations over states and coarse-graining operations over time does not exhibit commutativity. For instance, assuming $A_{m\times n}$ is a coarse-graining operation over states (merging $n$ states into $m$ states), ()×()($\cdot$)$\times$($\cdot$)( ⋅ ) × ( ⋅ ) is a coarse-graining operation over time (combining two time steps into one), the equation $A_{m\times n}(TPM_{n\times n})\times A_{m\times n}(TPM_{n\times n})=A_{m\times n}(TPM_{n\times n}\times TPM_{n\times n})$ does not always hold. This indicates that certain coarse-graining operations can result in a discrepancy between the evolution of macroscopic states and the coarse-grained states of the evolved microscopic systems. It implies the need for consistent constraints on coarse-graining strategies.
@@
That means solely maximizing EI may raise some problems, and further constraints must be added in the framework. We will discuss this problem in the section of 4.1.2.
3.8 Rosas Fernando’s Quantifying Causal Emergence
@@ 3.8.1 Basic idea
In Hoel’s framework, it is essential to find a coarse-graining strategy in order to determine the occurrence of causal emergence, and the outcome is influenced by the choice of the coarse-graining method. While it has been suggested by Hoel Hoel et al. (2013); Hoel (2017) that an optimal strategy can be identified by maximizing EI, certain issues have been raised as proposed by Dewhurst Dewhurst (2021). Third, EI is an global measure because it requires to intervene the input variable $X$ to be a uniform distribution on the whole domain Rosas et al. (2020). However, many regions are not be able to be observed from data. Consequently, there is a pressing need for an alternative theoretical framework of causal emergence that does not rely on a coarse-graining method.
In response, Fernando E. Rosas took an approach that did not require a coarse-graining strategy as prerequisite, attempting to break down excess entropy - the mutual information between a system’s past and future states - into non-overlapping parts to identify the information components most relevant to causal emergence. To accomplish this, he relied on the Partial Information Decomposition (PID) framework proposed by Williams and Beer, which provides a method for non-overlapping decomposition of joint mutual information Williams and Beer (2010). First, we will introduce Williams and Beer’s theoretic framework.
@@ 3.8.2 Partial Information Decomposition
The Partial Information Decomposition (PID) framework investigates the general informational relationship between source variables and a target variable. To make the problem description simpler without sacrificing generality, let us consider a system with two input variables ($X_{1}$, $X_{2}$) and one output variable ($Y$) as an example, as depicted in the Venn diagram below.
Figure 9: Venn diagram of PID. (a) Relationship of mutual information between Y and X1, X2. (b-d) Examples that joint mutual information is solely contributed by redundant, synergistic, and unique information. (e) Redundancy lattice representation of (a).
The mutual information between the target variable and individual source variables,$I(X_{1};Y)$ and $I(X_{2};Y)$, as well as the mutual information between the target variable and the joint source variable, $I(X_{1},X_{2};Y)$, exhibit a complex relationship and can not be intuitively converted to one another. Nonetheless, it’s an intuitive notion that the overlapped circular plates of $I(X_{1};Y)$ and $I(X_{2};Y)$ divide the oval region of $I(X_{1},X_{2};Y)$ into four adjacent non-overlapping regions, representing the three types of information components of $I(X_{1},X_{2};Y)$:
@@
If we can identify variable representations corresponding to these information components, the non-overlapping feature in the Venn diagram indicates that they must be independent of one another.
For a more intuitive understanding, let’s consider a few simple toy examples.
Case 1, $Y=X_{1}=X_{2}$: This is a scenario in which a single source variable can predict the target variable, and the addition of another source variable does not enhance the prediction of the target variable. In this case, we have $I(X_{1},X_{2};Y)=Red(X_{1},X_{2};Y)$, Fig 9(b).
Case 2, $Y=X_{1}\oplus X_{2},X_{1}\perp{X_{2}}$: In another scenario, neither of the source variables can predict the target variable individually, but together, they can predict the target variable synergistically. In this case, $I(X_{1},X_{2};Y)=Syn(X_{1},X_{2};Y)$, Fig 9(c).
Case 3, $Y=X_{1},X_{1}\perp{X_{2}}$: In this scenario, the target variable can be predicted by one of the source variables but not the other, which implies $I(X_{1},X_{2};Y)=Un(X_{1};Y|X_{2})$, Fig 9(d).
For general cases, Williams and Beer presented a method for calculating redundant information defined as $Red_{min}$. This method reflects the concept of redundancy, identifying it as the shared information across all sources, is equivalent to the minimum amount of information contributed by any single source:
@@
Fig 9(e) presents an alternative way of visualizing the PID framework, known as the redundancy lattice Williams and Beer (2010). In this representation, {12} denotes synergistic information, $\left\{1\right\}$ and $\left\{2\right\}$ denote unique information, and $\left\{1\right\}$$\left\{2\right\}$ denotes redundant information. In the subsequent sections, we will also see that the $\phi$ID framework employs the notation of the redundancy lattice.
@@ 3.8.3 Integrated Information Decomposition
The PID framework provides a useful framework for analyzing the non-overlapping information composition in a multivariate system. However, its application to causal analysis of dynamic systems is limited by the fact that it only allows for a single target variable. This limitation prevents the framework from fully capturing the transitions of multiple states across time steps. To address this challenge, Fernando E. Rosas etc have developed Integrated Information Decomposition ($\phi$ID) Mediano et al. (2019), which takes its name from Integrated Information Theory (IIT) Oizumi et al. (2014). This extension of PID provides a more comprehensive means for analyzing dynamic systems.
@@
To introduce Rosas’s framework clearly, we only consider a system with two variables. All the definitions and calculations can be generalized to the systems with more variables.
@@
The objective of the $\phi$ID framework is to decompose excess entropy into non-overlapping information components. In a two-variable Markovian system, the excess entropy is given by $E=I(X^{1}_{t},X^{2}_{t};X^{1}_{t+1},X^{2}_{t+1})$, where $X^{1}_{t}$ and $X^{2}_{t}$ represent the current states, and $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent the future states. From the causation perspective, $X^{1}_{t}$ and $X^{2}_{t}$ represent causes, while $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent effects.
@@
Rosas first noted that mathematically, the PID framework is symmetric with respect to source and target variables. There are two viewpoints for analyzing the above Markovian system with PID. One viewpoint takes $X^{1}_{t+1}$ and $X^{2}_{t+1}$ together as the target variable (aiming to decompose the "cause" to elements), while the other viewpoint takes $X^{1}_{t}$ and $X^{2}_{t}$ together as the target variable (aiming to decompose the "effect" to elements). The redundancy lattices of these perspectives are illustrated in Fig 10 (a) and (b), and respectively, are referred to as "forward PID" and " PID".
Figure 10: (a) Redundancy lattice of "forward PID", and its target variable $X^{1}_{t+1}X^{2}_{t+1}$. (b) Redundancy lattice of "backward PID" and its target variable $X^{1}_{t}X^{2}_{t}$.
@@
Next, Rosas introduced the $\phi$ID framework, which consolidates the forward PID and backward PID into a single framework. In this framework, the one-to-many relationship in PID is expanded to include many-to-many relationships. He built full connections between the elements of redundancy lattices of the forward PID and the backward PID. This approach generates 16 relations between the source and target, as depicted by the colored lines in Fig 11 (a). (Each color corresponds to a specific source element of $X_{t}$.) These relations are referred to as "$\phi$ID atoms" and can be represented as vertices in a lattice, as shown in Fig 11 (b). This more complex lattice is referred to as the double-redundancy lattice because it is a "product" of two redundancy lattices (of forward PID and backward PID).
@@
Figure 11: (a) Relationships between redundancy lattices of forward PID and backward PID. (b) Double-redundancy lattice, whose vertices correspond to the edges in (a)
@@
In the double-redundancy lattice, each vertex is defined as a "$\phi$ID atom," denoted as $I^{\alpha\rightarrow\beta}_{\partial},\alpha,\beta\in\mathcal{A}$, let $\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$. For instance, the $\phi$ID atom of the vertice ${\{12\}\rightarrow\{12\}}$ is demoted as $I^{\{12\}\rightarrow\{12\}}_{\partial}$. $\phi$ID atom represents the broken-down information transmitted from the cause elements to the effect elements, and their cumulative sum amounts to the excess entropy of the entire system. Consequently, the formal definition of Integrated Information Decomposition ($\phi$ID) is:
@@
The main advantage of $\phi$ID is that it can provide more fine-grained analysis of complex systems than PID by introducing the temporal dynamic. By decomposing the mutual information of the system at different times, $\phi$ID provides another approach for the quantitative study of Causal Emergence.

### 25 Feb 2024 v3

@@ 3.5.7 Other Applications
Once the method for quantifying causal emergence in complex systems is developed, it can be applied across various fields that possess abundant network data. The first type of network studied was biological networks.
@@
As previously discussed, biological networks are full of noise, which poses challenges in comprehending their internal operating principles. On one hand, such noise arises from inherent fluctuations within the system itself, whereas on the other hand, it can be introduced through measurement or observation processes. Consequently, Klein et al. Klein et al. (2022) further explored the relationship between noise, degeneracy, and certainty in biological networks and their specific meanings. For instance, in gene expression networks, highly deterministic relationships indicate that the expression of one gene almost invariably leads to the expression of another gene. Simultaneously, degeneracy is a prevalent phenomenon in the evolutionary processes of biological systems. Due to these two factors, it remains unclear at which scale biological systems should be analyzed to gain a deeper understanding of their functions.
@@
To address this, Klein et al. Klein et al. (2021) conducted an analysis of protein interaction networks across more than 1800 species. They employed EI as a measure for assessing the levels of noise and uncertainty in protein interactions. The findings revealed that the macro-scale network exhibited lower levels of noise and degeneracy. Additionally, the nodes within the macro-scale interaction group demonstrated greater resilience compared to the nodes that were not part of the macro-scale interaction group. Through robust analysis, the authors demonstrated that eukaryotes exhibited a stronger degree of causal emergence compared to archaea. Additionally, to address the ’deterministic paradox,’ the authors introduced the concepts of the neutral process and the selection process in biological evolution. The neutral process operates at the micro-scale, leveraging mutations to promote interactions and enhance species diversity. On the other hand, the selection process operates at the macro-scale, effectively eliminating noise that hampers system operation and efficiency. Therefore, in order to adapt to the demands of evolution, it becomes essential for evolved biological systems to function across multiple scales.
@@
Hoel et al. Hoel and Levin (2020) conducted further research on the causal emergence within biological systems. The authors elaborated that macro- and micro-systems exist widely in biological systems. For example, the micro-scale of a group of cells can involve potential ion channel changes, whereas the macro-scale corresponds to the membrane potential changes of cells. Furthermore, the authors utilized EI to analyze gene regulatory networks, aiming to identify the most informative models for controlling mammalian heart development. By quantifying the causal emergence within the largest component of the gene network of Saccharomyces cerevisia, it was revealed that informative macro-scale structures were prevalent across biological systems. Additionally, the authors emphasized the importance of evolutionary systems operating at multiple scales due to the significant advantages they offer. Natural selection requires variation between populations, and degradation observed in biological systems serves as a crucial factor for evolution. Degradation provides the conditions for the evolutionary process. However, organisms also need to maintain predicted consistency in phenotype, behavior, and structure to ensure survival and reproduction. Consequently, evolved systems need to operate at multiple scales, and the function of multi-scale systems is malleable in changing environments.
@@
Swain et al. Swain et al. (2022) conducted an investigation into the impact of ant colony interaction history on task assignment and task switching. They employed effective information to examine how noise information propagates among ants and explored the relationship between EI and the proportion of ants assigned to different tasks within ant colonies. The study revealed that the extent of historical interaction between ant colonies influenced task assignment. Additionally, the specific type of ant colony involved in an interaction determined the level of noise present within that interaction. For example, the interactions between foragers displayed significantly higher levels of noise when contrasted with interactions between nurses or cleaners. Furthermore, even when ants switched functional groups, the cohesion within ant colonies ensured the stability of the overall colony, and different functional ant colonies also played different roles in maintaining group cohesion.
@@
The EI indicator and causal emergence theoretic framework can also be applied to artificial systems. For example, Marrow et al. Marrow et al. (2020) quantified and monitored the changes in the causal structure of neural networks during training, in which EI was employed to evaluate the degree of causal influence of nodes and edges on downstream tasks at each layer. By observing the changes in EI, including determinism (sensitivity) and degeneracy, throughout model training, the generalization ability of the model could be determined, thus helping better understand and explain the working principle of neural networks.
@@
Varley et al. Varley (2020) attempted to apply the causal emergence framework to both discrete cellular automata and continuous Rossler systems. In the case of cellular automata systems, the authors select 88 unique rules corresponding to four types: static, periodic, chaotic, and complex. By considering each state as a node, where each state determines the subsequent state, a directed state transition graph is constructed. The analysis revealed that rules 1, 2, and 4 correspond to the strongest causal emergence of dynamics. Notably, the network constructed by these three rules exhibited a significant presence of star and spoke motifs. In addition, the authors also draw some quantitative conclusions, for example, among the 17 rules belonging to the third and fourth categories, 30% exist causal emergence, 70% show causal degradation, and the CE of cellular automata with the same rules remained relatively consistent across different sizes.
@@
Furthermore, the authors employed the OPN algorithm to transform the continuous system into a discrete-state transition graph for comparative analysis (see Section 3.5.5). They found that chaos dynamics demonstrated a correlation with low determinism, and the variations in the degeneracy and efficiency coefficients aligned with the changes observed in the determinism curve.
3.5.8 Critiques
@@
In the extensive literature on causality and emergence, Hoel’s theory has attracted attention for linking emergence and causality through interventionism, introducing the concept of causal emergence in a quantitative manner. However, Dewhurst Dewhurst (2021) provided a philosophical clarification of Hoel’s theory, arguing that it was epistemological rather than metaphysical. This suggests that Hoel’s macroscopic causality is merely a causal explanation based on information theory, rather than involving “genuinely novel causal powers”. This also raises concerns about the assumption of a uniform distribution, as there is no empirical basis to favor it over other distributions.
@@
The computation of Hoel’s effective information relies on two premises: (1) knowledge of the system’s microscopic dynamics, and (2) knowledge of the coarse-graining scheme. However, in practice, it is rare for both conditions to be simultaneously satisfied, especially in observational studies where both are unknown. This limitation hinders the practical applicability of Hoel’s theory.
@@
It has been pointed out that Hoel’s theory neglects the constraints on coarse-graining methods, and certain coarse-graining methods can lead to ambiguity Eberhardt and Lee (2022). Additionally, the combination of some coarse-graining operations over states and coarse-graining operations over time does not exhibit commutativity. For instance, by assuming $A_{m\times n}$ is a coarse-graining operation over states (merging $n$ states into $m$ states), and ()×()($\cdot$)$\times$($\cdot$)( ⋅ ) × ( ⋅ ) is a coarse-graining operation over time (combining two time steps into one), the equation $A_{m\times n}(TPM_{n\times n})\times A_{m\times n}(TPM_{n\times n})=A_{m\times n}(TPM_{n\times n}\times TPM_{n\times n})$ does not always hold. This indicates that certain coarse-graining operations can result in a discrepancy between the evolution of macroscopic states and the coarse-grained states of the evolved microscopic systems. It implies the need for consistent constraints on coarse-graining strategies.
@@
This means that solely maximizing EI may raise some problems, and further constraints must be added to the framework. We discuss this problem in Section 4.1.1.
3.6 Fernando E. Rosas’s Quantification of Causal Emergence
@@ 3.6.1 Basic Idea
In Hoel’s framework, it is essential to find a coarse-graining strategy in order to determine the occurrence of causal emergence, and the outcome is influenced by the choice of the coarse-graining method. Although it has been suggested by Hoel Hoel et al. (2013); Hoel (2017) that an optimal strategy can be identified by maximizing EI, certain issues have been raised by Dewhurst Dewhurst (2021). EI is a global measure because it requires the input variable $X$ to be a uniform distribution over the whole domain Rosas et al. (2020). However, many regions are not observable from the data. Consequently, there is a pressing need for an alternative theoretical framework for causal emergence that does not rely on a coarse-graining method.
In response, Fernando E. Rosas took an approach that did not require a coarse-graining strategy as a prerequisite, attempting to break down excess entropy—the mutual information between a system’s past and future states—into non-overlapping parts to identify the information components most relevant to causal emergence. To accomplish this, he relied on the partial information decomposition (PID) framework proposed by Williams and Beer, which provides a method for the non-overlapping decomposition of joint mutual information Williams and Beer (2010). Below, we introduce Williams and Beer’s theoretical framework.
@@ 3.6.2 Partial Information Decomposition
The partial information decomposition (PID) framework investigates the general informational relationship between source variables and a target variable. To simplify the problem description without sacrificing generality, let us consider a system with two input variables ($X_{1}$, $X_{2}$) and one output variable ($Y$) as an example, as depicted in the Venn diagram below.
The mutual information between the target variable and individual source variables, $I(X_{1};Y)$ and $I(X_{2};Y)$, as well as the mutual information between the target variable and the joint source variable, $I(X_{1},X_{2};Y)$, exhibits a complex relationship. Intuitively, one cannot be converted to another. Nonetheless, it is an intuitive notion that the overlapped circular plates of $I(X_{1};Y)$ and $I(X_{2};Y)$ divide the oval region of $I(X_{1},X_{2};Y)$ into four adjacent non-overlapping regions, representing the three types of information components of $I(X_{1},X_{2};Y)$:
@@
If we can identify variable representations corresponding to these information components, the non-overlapping feature in the Venn diagram indicates their independence from one another.
For a more intuitive understanding, let us consider a few simple toy examples.
Case 1, $Y=X_{1}=X_{2}$: This is a scenario in which a single source variable can predict the target variable, and the addition of another source variable does not enhance the prediction of the target variable. In this case, we have $I(X_{1},X_{2};Y)=Red(X_{1},X_{2};Y)$, as depicted in Figure 9b.
Case 2, $Y=X_{1}\oplus X_{2},X_{1}\perp{X_{2}}$: In another scenario, neither of the source variables can predict the target variable individually, but together, they can predict the target variable synergistically. In this case, $I(X_{1},X_{2};Y)=Syn(X_{1},X_{2};Y)$, as depicted in Figure 9c.
Case 3, $Y=X_{1},X_{1}\perp{X_{2}}$: In this scenario, the target variable can be predicted by one of the source variables but not the other, which implies $I(X_{1},X_{2};Y)=Un(X_{1};Y|X_{2})$, as depicted in Figure 9d.
Figure 9: Venn diagram of PID. (a) Relationship of mutual information between Y and X1, X2. (bd) Examples showing that joint mutual information is solely contributed by redundant, synergistic, and unique information. (e) Redundancy lattice representation of (a).
For general cases, Williams and Beer presented a method for calculating redundant information, defined as $Red_{min}$. This method reflects the concept of redundancy, identifying it as the shared information across all sources, which is equivalent to the minimum amount of information contributed by any single source:
@@
Figure 9e presents an alternative way of visualizing the PID framework, known as the redundancy lattice Williams and Beer (2010). In this representation, {12} denotes synergistic information, $\left\{1\right\}$ and $\left\{2\right\}$ denote unique information, and $\left\{1\right\}$$\left\{2\right\}$ denotes redundant information. In the subsequent sections, we demonstrate that the $\phi$ID framework employs the notation of the redundancy lattice.
@@ 3.6.3 Integrated Information Decomposition
The PID framework provides a useful framework for analyzing the non-overlapping information composition in a multivariate system. However, its application to causal analysis of dynamic systems is limited by the fact that it only allows for a single target variable. This limitation prevents the framework from fully capturing the transitions of multiple states across time steps. To address this challenge, Fernando E. Rosas developed Integrated Information Decomposition ($\phi$ID) Mediano et al. (2019), which takes its name from Integrated Information Theory (IIT) Oizumi et al. (2014). This extension of PID provides a more comprehensive method for analyzing dynamic systems.
@@
To introduce Rosas’s framework clearly, we consider a system with only two variables. All the definitions and calculations can be generalized to systems with more variables.
@@
The objective of the $\phi$ID framework is to decompose excess entropy into non-overlapping information components. In a two-variable Markovian system, excess entropy is given by $E=I(X^{1}_{t},X^{2}_{t};X^{1}_{t+1},X^{2}_{t+1})$, where $X^{1}_{t}$ and $X^{2}_{t}$ represent the current states, and $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent the future states. From a causation perspective, $X^{1}_{t}$ and $X^{2}_{t}$ represent causes, whereas $X^{1}_{t+1}$ and $X^{2}_{t+1}$ represent effects.
@@
Rosas first noted that mathematically, the PID framework is symmetric with respect to the source and target variables. There are two viewpoints for analyzing the above Markovian system with PID. One viewpoint takes $X^{1}_{t+1}$ and $X^{2}_{t+1}$ together as the target variable (aiming to decompose the “cause” to elements), whereas the other viewpoint takes $X^{1}_{t}$ and $X^{2}_{t}$ together as the target variable (aiming to decompose the “effect” to elements). The redundancy lattices of these perspectives are illustrated in Figure 10a,b, and respectively, are referred to as “forward PID” and “ PID”.
Figure 10: (a) Redundancy lattice of “forward PID” and its target variable $X^{1}_{t+1}X^{2}_{t+1}$. (b) Redundancy lattice of “backward PID” and its target variable $X^{1}_{t}X^{2}_{t}$.
@@
Next, Rosas introduced the $\phi$ID framework, consolidating forward PID and backward PID into a single framework. In this framework, the one-to-many relationship in PID was expanded to include many-to-many relationships. He built full connections between the elements of the redundancy lattices of forward PID and backward PID. This approach generates 16 relations between the source and target, as depicted by the colored lines in Figure 11a (each color corresponds to a specific source element of $X_{t}$). These relations are referred to as “$\phi$ID atoms” and can be represented as vertices in a lattice, as shown in Figure 11b. This more complex lattice is referred to as the double-redundancy lattice because it is a “product” of two redundancy lattices (of forward PID and backward PID).
@@
Figure 11: (a) Relationships between redundancy lattices of forward PID and backward PID. (b) Double-redundancy lattice, whose vertices correspond to the edges in (a). Distinct colors of dots and lines are employed to distinguish the source elements of $X_{t}$.
@@
In the double-redundancy lattice, each vertex is defined as a “$\phi$ID atom”, denoted as $I^{\alpha\rightarrow\beta}_{\partial},\alpha,\beta\in\mathcal{A}$, where $\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$. For instance, the $\phi$ID atom of the vertice ${\{12\}\rightarrow\{12\}}$ is denoted as $I^{\{12\}\rightarrow\{12\}}_{\partial}$. The $\phi$ID atom represents the broken-down information transmitted from the cause elements to the effect elements, and their cumulative sum amounts to the excess entropy of the entire system. Consequently, the formal definition of Integrated Information Decomposition ($\phi$ID) is:
@@
The main advantage of $\phi$ID is that it can provide a more fine-grained analysis of complex systems compared to PID by introducing the temporal dynamic. By decomposing the mutual information of the system at different times, $\phi$ID provides another approach for the quantitative study of causal emergence.